THE SEARCH FOR A DOUBLE-STUFFED URANIUM METALLOCENE

\underline{J. L. SONNENBERG}, J. ZHOU, H. B. SCHLEGEL, Department of Chemistry, Wayne State University, Detroit, MI-48202.

The discovery of Zn(I)_{2}(η^{5}-C_{5}Me_{5})^{2} has fundamentally changed the definition of metallocene by introducing dimetallic units to the middle of the classic sandwich complex. Since actinide metallocenes employing COT (COT = [C_{8}H_{8}]^{2−}) rings are well known, it occurred to us that uranium dimers may be stabilized by this new bonding motif. Scalar-relativistic density functional theory calculations have been employed to investigate the isomers of [U(n)_{2}(COT)_{2}]^{2n−4} complexes where n = 2 or 3. The IR and NMR vibrational signatures of these complexes is presented to facilitate experimental detection.